OpenStudy (anonymous):

simplify square root: 3√ 2 + 6√ 2 - 4√ 3 I am SOOOOO confused

OpenStudy (anonymous):

3√ 2 + 6√ 2 - 4√ 3 =9√ 2 - 4√ 3

OpenStudy (anonymous):

can i ask how you did that?

OpenStudy (anonymous):

Yeh sure

OpenStudy (anonymous):

Check out the radicals which have same radical and they can be added

OpenStudy (anonymous):

3√ 2 + 6√ 2 - 4√ 3 = (3+6)√ 2-4√ 3=9√ 2-4√ 3

OpenStudy (anonymous):

since √ 2 & √ 3 are two different numbers so they can not be added or subtracted

OpenStudy (anonymous):

ok